OpenStudy (anonymous):

Use the following equation: log2x + log2(x – 7) = 3 To solve the equation, first: Select one: a. Use the quotient property of logarithms to condense the equation. b. Use the product property of logarithms to condense the equation. c. Use the power property of logarithms to condense the equation. d. Either B or C

OpenStudy (anonymous):

@IMStuck

myininaya (myininaya):

Here are some properties of log you might find useful: log(a)-log(b)=log(a/b) <--quotient rule because it has a quotient inside the log log(a)+log(b)=log(ab) <--product rule because it has a product inside the log b*log(a)=log(a^b) <--power rule because there is a power in the log) Which rule do you think I have listed applies here? Compare your left hand side to the left hand sides of the properties I just wrote.

OpenStudy (anonymous):

@ParthKohli